# Convergence and Generalization of Anti-Regularization for Parametric Models

### v1

@@ A.2 Basic Spectral Facts
Also, if $\lambda<\sigma_{\min}(A)$, then $\|(A-\lambda I)^{-1}\|_{2}=\frac{1}{\sigma_{\min}(A)-\lambda}$.
Lemma 9 (Trace Monotonicity and Differentiation).
Given the eigendecomposition $A=U\operatorname{diag}(\sigma_{j})U^{\top}$ and scalar $\lambda$,
@@
More generally, for invertible $M(\lambda)$,
$\frac{\partial}{\partial\lambda}M^{-1}=-M^{-1}\left(\frac{\partial M}{\partial\lambda}\right)M^{-1}.$
Lemma 10 (PSD Ordering and Trace).
@@
Lemma 11 (Sherman–Morrison–Woodbury Formula).
For invertible $A$ and compatible dimensions of $U,C,V$,
@@
This is especially useful when $A=X^{\top}X-n\lambda I$, enabling efficient updates of $A^{-1}$.
Lemma 12 (Weyl’s Inequality (Simplified)).
@@
Remark 1 (Spectral Expression for Degrees of Freedom).
@@
A.3 Optimism Identity Background
This section summarizes the derivation of the following identity under the linear smoother $\hat{f}_{\lambda}=S_{\lambda}y$:
$\mathbb{E}\!\left[\mathcal{R}(\hat{f}_{\lambda})\right]=\mathbb{E}\!\left[\widehat{\mathcal{R}}_{S}(\hat{f}_{\lambda})\right]+\frac{2\tau^{2}}{n}\operatorname{tr}(S_{\lambda}),$
which belongs to the same family of results as Stein’s Unbiased Risk Estimate (SURE) and the optimism identity in prediction.
@@ A.3 Optimism Identity Background
Assume the generative model $y=X\theta^{\star}+\varepsilon$, where $\varepsilon\sim(0,\tau^{2}I)$, and define the smoother $\hat{f}_{\lambda}=S_{\lambda}y$. Using the definitions of $\widehat{\mathcal{R}}_{S}$ and $\mathcal{R}$ from §A.1, we have
@@
Start from the decomposition: $\hat{f}_{\lambda}=S_{\lambda}(X\theta^{\star}+\varepsilon)=S_{\lambda}X\theta^{\star}+S_{\lambda}\varepsilon$. For new responses $y^{\text{new}}=X\theta^{\star}+\varepsilon^{\prime}$, we compute
$\frac{1}{n}\mathbb{E}_{\varepsilon^{\prime}}\!\left[\|y^{\text{new}}-\hat{f}_{\lambda}\|^{2}\right]=\frac{1}{n}\|X\theta^{\star}-\hat{f}_{\lambda}\|^{2}+\tau^{2}.$
On the other hand,
@@
Taking expectation over $\varepsilon$, the cross term vanishes, and the last term becomes
$\frac{1}{n}\mathbb{E}\!\left[\operatorname{tr}\left((I-S_{\lambda})^{2}\varepsilon\varepsilon^{\top}\right)\right]=\frac{\tau^{2}}{n}\operatorname{tr}\left((I-S_{\lambda})^{2}\right).$
By definition, $\widehat{\mathcal{R}}_{S}=\tfrac{1}{n}\|y-\hat{f}_{\lambda}\|^{2}-\tau^{2}$, so we obtain
$\mathbb{E}\!\left[\mathcal{R}\right]=\mathbb{E}\!\left[\widehat{\mathcal{R}}_{S}\right]+\frac{\tau^{2}}{n}\left\{\operatorname{tr}((I-S_{\lambda})^{2})-\operatorname{tr}(I)\right\}.$
For linear smoothers, the identity $\operatorname{tr}((I-S_{\lambda})^{2})-\operatorname{tr}(I)=-2\,\operatorname{tr}(S_{\lambda})$ holds, which completes the proof. ∎
Remark 2 (Degrees of Freedom Interpretation).
In the above identity, $\operatorname{tr}(S_{\lambda})$ acts as the model degrees of freedom. Thus, strategies that control $\operatorname{tr}(S_{\lambda})$ to remain constant or decay with sample size offer a principled way to manage the bias–variance tradeoff.
Appendix B Regression: Unique Convergence in the Safe Zone
@@ B.1 Proof of Strong Convexity
Let the design matrix be $X\in\mathbb{R}^{n\times p}$ and $\hat{\Sigma}=\tfrac{1}{n}X^{\top}X$. For the regression loss $\ell(\theta;x,y)=\tfrac{1}{2}(y-x^{\top}\theta)^{2}$ and reward $R(\theta)=\tfrac{1}{2}\|\theta\|_{2}^{2}$, the objective
$\hat{F}_{\lambda}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\tfrac{1}{2}(y_{i}-x_{i}^{\top}\theta)^{2}\;-\;\frac{\lambda}{2}\|\theta\|_{2}^{2}$
@@
We compute $\nabla^{2}\hat{F}_{\lambda}=\hat{\Sigma}-\lambda I$. By spectral bounds of symmetric matrices, we have $\sigma_{\min}(H_{\lambda})=\sigma_{\min}(\hat{\Sigma})-\lambda$. By assumption, this quantity is positive, i.e., $H_{\lambda}\succ 0$, so $\hat{F}_{\lambda}$ is $\mu$-strongly convex and has a unique minimizer. ∎
@@
$\hat{F}_{\lambda}(tu)=\frac{1}{2}t^{2}(\sigma_{\min}(\hat{\Sigma})-\lambda)-\frac{t}{n}y^{\top}Xu+\text{const},$
and the coefficient of $t^{2}$ is negative, so $\hat{F}_{\lambda}(tu)\to$-$\infty$ as $t\to\infty$.
Remark 4 (Extension to Weighted Rewards).
If the reward term is generalized to $R(\theta)=\tfrac{1}{2}\theta^{\top}W\theta$ with $W\succeq 0$, then the Hessian becomes $H_{\lambda}=\hat{\Sigma}-\lambda W$. The safe condition is then given in terms of the generalized eigenvalue problem as $\lambda<\inf_{\|v\|=1}\tfrac{v^{\top}\hat{\Sigma}v}{v^{\top}Wv}$. When $W=I$, this reduces to the original theorem.
B.2 Closed-Form Solution and Gradient Descent Convergence
@@
Under the assumptions of the theorem, the first-order optimality condition $\nabla\hat{F}_{\lambda}(\theta)=0$ becomes
@@
The solution is unique and exists since $H_{\lambda}\succ 0$.
@@
We compute $\nabla\hat{F}_{\lambda}(\theta)=\hat{\Sigma}\theta-\tfrac{1}{n}X^{\top}y-\lambda\theta$. Setting this to zero yields a linear equation, and since $H_{\lambda}$ is invertible, the solution follows. ∎
@@
Consider standard gradient descent on the objective $\hat{F}_{\lambda}$: $\theta_{t+1}=\theta_{t}-\eta\nabla\hat{F}_{\lambda}(\theta_{t})$. Let $L=\sigma_{\max}(H_{\lambda})=\sigma_{\max}(\hat{\Sigma})-\lambda$, and $\mu=\sigma_{\min}(H_{\lambda})=\sigma_{\min}(\hat{\Sigma})-\lambda$. If $0<\eta<\tfrac{2}{L}$, then
$\|\theta_{t}-\hat{\theta}_{\lambda}\|_{2}\;\leq\;\rho^{t}\|\theta_{0}-\hat{\theta}_{\lambda}\|_{2},\qquad\rho=\max_{j}|1-\eta\sigma_{j}(H_{\lambda})|<1,$
i.e., the convergence is linear. The optimal fixed step size is $\eta^{\star}=\tfrac{2}{L+\mu}$, yielding rate $\rho^{\star}=\tfrac{\kappa-1}{\kappa+1}$ with condition number $\kappa=\tfrac{L}{\mu}$.
@@
Since the objective is quadratic, the error $e_{t}=\theta_{t}-\hat{\theta}_{\lambda}$ evolves as $e_{t+1}=(I-\eta H_{\lambda})e_{t}$, implying $\|e_{t+1}\|_{2}\leq\|I-\eta H_{\lambda}\|_{2}\|e_{t}\|_{2}$. By spectral decomposition, $\|I-\eta H_{\lambda}\|_{2}=\max_{j}|1-\eta\sigma_{j}(H_{\lambda})|$. For $\eta\in(0,2/L)$, this quantity is less than 1, guaranteeing linear convergence. The expressions for $\eta^{\star}$ and $\rho^{\star}$ follow from classical optimization theory. ∎
Remark 5 (Step Size and Numerical Stability).
In practice, upper bounds on $L$ are often overestimated, so a conservative step size such as $\eta\leq\tfrac{1}{L}$ is used, possibly with outer line search. When using a mini-batch variant, due to variance in $H_{\lambda}$, reducing $\eta$ further is often advisable for stability.
B.3 Comments and Counterexamples Near the Boundary
@@
$\left\|(X^{\top}X-n\lambda I)^{-1}\right\|_{2}=\tfrac{1}{n(\sigma_{\min}(\hat{\Sigma})-\lambda)}\to\infty.$
Practical Safeguards
@@
Let $\hat{\Sigma}=\operatorname{diag}(1,\delta)$, $0<\delta<1$, and $\lambda=\tfrac{1+\delta}{2}$. Then,
$H_{\lambda}=\operatorname{diag}\left(\tfrac{1-\delta}{2},-\tfrac{1-\delta}{2}\right)$
has a negative eigenvalue. Along the direction $u=(0,1)^{\top}$, the objective $\hat{F}_{\lambda}(tu)\to$-$\infty$ as $t\to\infty$, implying the objective is unbounded below if $\lambda$ exceeds the spectral bound.
@@
Instability near the boundary can be mitigated through: (i) Reducing $\lambda$ further or applying a decay schedule $|\lambda(n)|\propto n^{-\alpha}$ as sample size increases; (ii) Temporarily combining a trust region constraint $\|\theta\|\leq B$ to prevent numerical blow-up; (iii) Aligning the coordinate system $X=U\Sigma^{1/2}V^{\top}$ and selectively damping the small eigenspace directions.
@@ Appendix C Degrees of Freedom and Test Risk Improvement
C.1 Spectral Derivation of Degrees of Freedom
Proposition 17 (Spectral Expression for Degrees of Freedom).
Let the eigenvalues of $\hat{\Sigma}=\tfrac{1}{n}X^{\top}X$ be $\{\sigma_{j}\}_{j=1}^{r}$, where $r=\operatorname{rank}(X)$. Then, for the linear smoother
$S_{\lambda}=X\left(X^{\top}X-n\lambda I\right)^{-1}X^{\top}\in\mathbb{R}^{n\times n},$
@@ C.1 Spectral Derivation of Degrees of Freedom
Using the cyclic property of the trace:
$\operatorname{tr}(S_{\lambda})=\operatorname{tr}\left(\left(X^{\top}X-n\lambda I\right)^{-1}X^{\top}X\right).$
Given $\hat{\Sigma}=V\operatorname{diag}(\sigma_{j})V^{\top}$, we write $X^{\top}X=nV\operatorname{diag}(\sigma_{j})V^{\top}$, and thus:
$\left(X^{\top}X-n\lambda I\right)^{-1}=\left(nV\operatorname{diag}(\sigma_{j}-\lambda)V^{\top}\right)^{-1}=\tfrac{1}{n}V\operatorname{diag}\left(\tfrac{1}{\sigma_{j}-\lambda}\right)V^{\top}.$
So we get:
$\operatorname{tr}(S_{\lambda})=\operatorname{tr}\left(V\operatorname{diag}\left(\tfrac{1}{\sigma_{j}-\lambda}\right)V^{\top}\cdot V\operatorname{diag}(\sigma_{j})V^{\top}\right)=\sum_{j=1}^{r}\frac{\sigma_{j}}{\sigma_{j}-\lambda}.$
Eigenvalues that are zero do not contribute to the trace. ∎
@@
With $X=U\Sigma^{1/2}V^{\top}$ and row-orthogonal $U$, the smoother becomes
$S_{\lambda}=U\operatorname{diag}\left(\tfrac{\sigma_{j}}{\sigma_{j}-\lambda}\right)U^{\top},$
and each mode contributes a smoothing coefficient $h_{j}(\lambda)=\tfrac{\sigma_{j}}{\sigma_{j}-\lambda}$ to the total degrees of freedom.
C.2 Proof of Test Risk Improvement Theorem
Theorem 18 (Sufficient Condition for Test Risk Improvement — Restatement).
Let
$\widehat{\mathcal{R}}_{S}(\hat{f})=\tfrac{1}{n}\|y-\hat{f}\|_{2}^{2}-\tau^{2},\quad\mathcal{R}(\hat{f})=\tfrac{1}{n}\mathbb{E}_{\varepsilon^{\prime}}\left[\|y^{\text{new}}-\hat{f}\|_{2}^{2}\right],$
and define
$\Delta\widehat{\mathcal{R}}_{S}:=\widehat{\mathcal{R}}_{S}(\hat{f}_{0})-\widehat{\mathcal{R}}_{S}(\hat{f}_{\lambda})\geq 0,\qquad\Delta\operatorname{tr}:=\operatorname{tr}(S_{\lambda})-\operatorname{tr}(S_{0})\geq 0.$
Then, from the optimism identity
we obtain the implication:

### v2

@@ A.2 Basic Facts on Spectral Matrices
Moreover, if $\lambda<\sigma_{\min}(A)$, then $\|(A-\lambda I)^{-1}\|_{2}=\tfrac{1}{\sigma_{\min}(A)-\lambda}$.
Lemma 10 (Trace monotonicity and differentiation rule).
For the diagonalization $A=U\operatorname{diag}(\sigma_{j})U^{\top}$ and scalar $\lambda$,
@@
In general, for invertible $M(\lambda)$,
$\frac{\partial}{\partial\lambda}M^{-1}=-M^{-1}\Big{(}\frac{\partial M}{\partial\lambda}\Big{)}M^{-1}.$
Lemma 11 (PSD order and trace).
@@
Lemma 12 (Sherman–Morrison–Woodbury).
For invertible $A$ and conformable $U,C,V$,
@@
In particular, when $A=X^{\top}X-n\lambda I$, this provides a useful update formula for $A^{-1}$.
Lemma 13 (Weyl’s inequality (sketch)).
@@
Remark 4 (Spectral representation of degrees of freedom).
@@
A.3 Background on the Optimism Identity
This section summarizes the derivation of the identity
$\mathbb{E}\!\left[\mathcal{R}(\hat{f}_{\lambda})\right]=\mathbb{E}\!\left[\widehat{\mathcal{R}}_{S}(\hat{f}_{\lambda})\right]+\frac{2\tau^{2}}{n}\operatorname{tr}(S_{\lambda})$
under the linear smoother $\hat{f}_{\lambda}=S_{\lambda}y$. This result belongs to the same family as Stein’s unbiased risk estimate (SURE) and the optimism identity for prediction risk.
@@ A.3 Background on the Optimism Identity
Consider the generative model $y=X\theta^{\star}+\varepsilon$ with $\varepsilon\sim(0,\tau^{2}I)$, and let $\hat{f}_{\lambda}=S_{\lambda}y$. With $\widehat{\mathcal{R}}_{S},\ \mathcal{R}$ defined as in §A.1,
@@
We expand $\hat{f}_{\lambda}=S_{\lambda}(X\theta^{\star}+\varepsilon)=S_{\lambda}X\theta^{\star}+S_{\lambda}\varepsilon$. For a new response $y^{\text{new}}=X\theta^{\star}+\varepsilon^{\prime}$,
$\frac{1}{n}\mathbb{E}_{\varepsilon^{\prime}}\!\big{[}\|y^{\text{new}}-\hat{f}_{\lambda}\|^{2}\big{]}=\frac{1}{n}\|X\theta^{\star}-\hat{f}_{\lambda}\|^{2}+\tau^{2}.$
Meanwhile,
@@
Taking expectation over $\varepsilon$, the cross term vanishes, and the last term becomes $\tfrac{1}{n}\mathbb{E}\big{[}\operatorname{tr}\big{(}(I-S_{\lambda})^{2}\varepsilon\varepsilon^{\top}\big{)}\big{]}=\tfrac{\tau^{2}}{n}\operatorname{tr}\big{(}(I-S_{\lambda})^{2}\big{)}$. By definition, $\widehat{\mathcal{R}}_{S}=\tfrac{1}{n}\|y-\hat{f}_{\lambda}\|^{2}-\tau^{2}$. Rearranging,
$\mathbb{E}\!\left[\mathcal{R}\right]=\mathbb{E}\!\left[\widehat{\mathcal{R}}_{S}\right]+\frac{\tau^{2}}{n}\Big{\{}\operatorname{tr}\big{(}(I-S_{\lambda})^{2}\big{)}-\operatorname{tr}(I)\Big{\}}.$
For linear smoothers, $\operatorname{tr}\big{(}(I-S_{\lambda})^{2}\big{)}-\operatorname{tr}(I)=-2\,\operatorname{tr}(S_{\lambda})$, yielding the stated identity. ∎
Remark 5 (Degrees-of-freedom interpretation).
In this identity, $\operatorname{tr}(S_{\lambda})$ acts as the model’s degrees of freedom. Thus, controlling $\lambda$ so that $\operatorname{tr}(S_{\lambda})$ is maintained at a constant proportion per sample, or decreases appropriately, provides a quantitative strategy to balance overfitting and underfitting.
Appendix B Regression: Safe Region, Unique Solution, and Convergence
@@ B.1 Proof of Strong Convexity
Let the design matrix be $X\in\mathbb{R}^{n\times p}$ and define $\hat{\Sigma}=\tfrac{1}{n}X^{\top}X$. Consider the regression loss $\ell(\theta;x,y)=\tfrac{1}{2}(y-x^{\top}\theta)^{2}$ and reward term $R(\theta)=\tfrac{1}{2}\|\theta\|_{2}^{2}$. The objective
$\hat{F}_{\lambda}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\tfrac{1}{2}\big{(}y_{i}-x_{i}^{\top}\theta\big{)}^{2}\;-\;\frac{\lambda}{2}\|\theta\|_{2}^{2}$
@@
Since $\nabla^{2}\hat{F}_{\lambda}=\hat{\Sigma}-\lambda I$, the minimum eigenvalue satisfies $\sigma_{\min}(H_{\lambda})=\sigma_{\min}(\hat{\Sigma})-\lambda$. By assumption, $\sigma_{\min}(H_{\lambda})>0$, hence $H_{\lambda}\succ 0$. Therefore, $\hat{F}_{\lambda}$ is $\mu$-strongly convex, and the uniqueness of the minimizer follows. ∎
@@
$\hat{F}_{\lambda}(tu)=\frac{1}{2}t^{2}\big{(}\sigma_{\min}(\hat{\Sigma})-\lambda\big{)}-\frac{t}{n}y^{\top}Xu+\text{const},$
whose quadratic coefficient is negative, so $\hat{F}_{\lambda}(tu)\to$-$\infty$ as $t\to\infty$.
Remark 7 (Extension to weighted reward).
If the reward is replaced by $R(\theta)=\tfrac{1}{2}\theta^{\top}W\theta$ with $W\succeq 0$, the Hessian becomes $H_{\lambda}=\hat{\Sigma}-\lambda W$. The safe condition is then given in terms of a generalized eigenvalue problem as
$\lambda<\inf_{\|v\|=1}\frac{v^{\top}\hat{\Sigma}v}{v^{\top}Wv}.$
When $W=I$, this reduces to the condition in the theorem above.
B.2 Closed-form Solution and Convergence of Gradient Descent
@@
Under the assumptions of the theorem, the first-order optimality condition $\nabla\hat{F}_{\lambda}(\theta)=0$ yields
@@
The solution exists and is unique since $H_{\lambda}\succ 0$.
@@
We have $\nabla\hat{F}_{\lambda}(\theta)=\hat{\Sigma}\theta-\tfrac{1}{n}X^{\top}y-\lambda\theta$. Setting this to zero gives a linear system, and because $H_{\lambda}$ is invertible we obtain the stated solution. ∎
@@
Consider standard gradient descent (GD) on $\hat{F}_{\lambda}$: $\theta_{t+1}=\theta_{t}-\eta\nabla\hat{F}_{\lambda}(\theta_{t})$. Let $L=\sigma_{\max}(H_{\lambda})=\sigma_{\max}(\hat{\Sigma})-\lambda$ and $\mu=\sigma_{\min}(H_{\lambda})=\sigma_{\min}(\hat{\Sigma})-\lambda$. If $0<\eta<\tfrac{2}{L}$, then
$\|\theta_{t}-\hat{\theta}_{\lambda}\|_{2}\;\leq\;\rho^{t}\,\|\theta_{0}-\hat{\theta}_{\lambda}\|_{2},\qquad\rho=\max_{j}\bigl{|}1-\eta\sigma_{j}(H_{\lambda})\bigr{|}<1,$
i.e., GD converges linearly. The optimal fixed stepsize is $\eta^{\star}=\tfrac{2}{L+\mu}$, for which $\rho^{\star}=\tfrac{\kappa-1}{\kappa+1}$ with condition number $\kappa=\tfrac{L}{\mu}$.
@@
Since the objective is quadratic, the error $e_{t}=\theta_{t}-\hat{\theta}_{\lambda}$ evolves as $e_{t+1}=(I-\eta H_{\lambda})e_{t}$. Hence $\|e_{t+1}\|_{2}\leq\|I-\eta H_{\lambda}\|_{2}\,\|e_{t}\|_{2}$, and by spectral decomposition $\|I-\eta H_{\lambda}\|_{2}=\max_{j}|1-\eta\sigma_{j}(H_{\lambda})|$. When $\eta\in(0,2/L)$, this quantity is strictly less than 1, implying linear convergence. The expressions for $\eta^{\star}$ and $\rho^{\star}$ are standard results in optimization. ∎
Remark 8 (Stepsize selection and numerical stability).
In practice, upper bounds on $L$ are often overestimated; it is therefore prudent to use $\eta\leq\tfrac{1}{L}$ or to employ an outer line search. When using mini-batches, additional reduction of $\eta$ is advisable due to variance in $H_{\lambda}$.
B.3 Remarks Near the Boundary and Counterexamples
@@
Practical guardrails
@@
Let $\hat{\Sigma}=\operatorname{diag}(1,\delta)$ with $0<\delta<1$, and set $\lambda=\tfrac{1+\delta}{2}$. Then $H_{\lambda}=\operatorname{diag}\big{(}\tfrac{1-\delta}{2},-\tfrac{1-\delta}{2}\big{)}$, which has a negative eigenvalue. Along the direction $u=(0,1)^{\top}$, setting $\theta=tu$ leads to $\hat{F}_{\lambda}(tu)\to$-$\infty$. Thus, once the boundary is exceeded, the objective becomes unbounded below.
@@
Instability near the boundary can be mitigated in several ways: (i) reduce $\lambda$ further or apply a decay schedule $|\lambda(n)|\propto n^{-\alpha}$ as the sample size grows; (ii) temporarily impose a trust-region constraint $\|\theta\|\leq B$ to prevent numerical blow-up; (iii) apply spectral alignment by decomposing $X=U\Sigma^{1/2}V^{\top}$ and attenuating only within the subspace of small eigenvalues.
@@ Appendix C Degrees of Freedom and Test Risk Improvement
C.1 Derivation of Degrees of Freedom for Linear Smoothers
Proposition 18 (Spectral representation of degrees of freedom).
Let $\{\sigma_{j}\}_{j=1}^{r}$ be the eigenvalues of $\hat{\Sigma}=\tfrac{1}{n}X^{\top}X$, where $r=\operatorname{rank}(X)$. For the linear smoother
$S_{\lambda}\;=\;X\big{(}X^{\top}X-n\lambda I\big{)}^{-1}X^{\top}\in\mathbb{R}^{n\times n},$
@@ C.1 Derivation of Degrees of Freedom for Linear Smoothers
By cyclicity of the trace,
$\operatorname{tr}(S_{\lambda})=\operatorname{tr}\!\Big{(}\big{(}X^{\top}X-n\lambda I\big{)}^{-1}X^{\top}X\Big{)}.$
With the eigen-decomposition $\hat{\Sigma}=V\operatorname{diag}(\sigma_{j})V^{\top}$, we have $X^{\top}X=nV\operatorname{diag}(\sigma_{j})V^{\top}$, and
$\big{(}X^{\top}X-n\lambda I\big{)}^{-1}=\big{(}nV\operatorname{diag}(\sigma_{j}-\lambda)V^{\top}\big{)}^{-1}=\tfrac{1}{n}\,V\operatorname{diag}\!\Big{(}\tfrac{1}{\sigma_{j}-\lambda}\Big{)}V^{\top}.$
Thus,
$\operatorname{tr}(S_{\lambda})=\operatorname{tr}\!\Big{(}V\operatorname{diag}\!\Big{(}\tfrac{1}{\sigma_{j}-\lambda}\Big{)}V^{\top}\cdot V\operatorname{diag}(\sigma_{j})V^{\top}\Big{)}=\sum_{j=1}^{r}\frac{\sigma_{j}}{\sigma_{j}-\lambda}.$
Eigenvalues equal to zero do not contribute to the sum. ∎
@@
With the left-orthogonal decomposition $X=U\Sigma^{1/2}V^{\top}$, we have $S_{\lambda}=U\,\operatorname{diag}\!\big{(}\tfrac{\sigma_{j}}{\sigma_{j}-\lambda}\big{)}\,U^{\top}$. Thus, each mode contributes a smoothing coefficient $h_{j}(\lambda)=\tfrac{\sigma_{j}}{\sigma_{j}-\lambda}$ to the degrees of freedom.
Remark 10 (Scope of the degrees-of-freedom formula).
[…]
C.2 Proof of the Test Risk Improvement Theorem
Theorem 19 (Sufficient condition for test risk improvement — restated).
Let $\widehat{\mathcal{R}}_{S}(\hat{f})=\tfrac{1}{n}\|y-\hat{f}\|_{2}^{2}-\tau^{2}$ and $\mathcal{R}(\hat{f})=\tfrac{1}{n}\mathbb{E}_{\varepsilon^{\prime}}\big{[}\|y^{\text{new}}-\hat{f}\|_{2}^{2}\big{]}$. Define $\Delta\widehat{\mathcal{R}}_{S}=\widehat{\mathcal{R}}_{S}(\hat{f}_{0})-\widehat{\mathcal{R}}_{S}(\hat{f}_{\lambda})\geq 0$ and $\Delta\operatorname{tr}=\operatorname{tr}(S_{\lambda})-\operatorname{tr}(S_{0})\geq 0$. From the optimism identity $\mathbb{E}[\mathcal{R}(\hat{f}_{\lambda})]=\mathbb{E}[\widehat{\mathcal{R}}_{S}(\hat{f}_{\lambda})]+\tfrac{2\tau^{2}}{n}\operatorname{tr}(S_{\lambda})$, it follows that